# Double Robust high dimensional alpha test for linear factor pricing model

### v1

@@ A.1 Proof of Theorem 1
$+\hat{r}_{t}^{-2}\hat{\boldsymbol{\theta}}^{\top}\hat{{\bf D}}^{-1}\boldsymbol{\varepsilon}^{\top}{\bf F}({\bf F}^{\top}{\bf F})^{-1}{\boldsymbol{f}}_{t}+\delta_{1t})\hat{{\bf U}}_{t}+\frac{1}{T}\sum_{t=1}^{T}(\hat{r}_{t}^{-1}\hat{{\bf U}}_{t}\hat{{\bf U}}_{t}^{\top}\hat{{\bf D}}^{-1/2}\hat{\boldsymbol{\theta}}+\hat{r}_{t}^{-1}\hat{{\bf U}}_{t}\hat{{\bf U}}_{t}^{\top}\hat{{\bf D}}^{-1/2}\boldsymbol{\varepsilon}^{\top}{\boldsymbol{V}}_{t})$
@@
$+\hat{r}^{-1}_{t}\hat{{\bf U}}^{\top}_{t}\hat{{\bf D}}^{1/2}\boldsymbol{\varepsilon}^{\top}{\bf F}({\bf F}^{\top}{\bf F})^{-1}{\boldsymbol{f}}_{t}+\hat{r}^{-2}_{t}\hat{\boldsymbol{\theta}}^{\top}\hat{{\bf D}}^{-1}\boldsymbol{\varepsilon}^{\top}{\bf F}({\bf F}^{\top}{\bf F})^{-1}{\boldsymbol{f}}_{t}$
@@
Firstly, we focus on the first term on the left side of (5). Due to Lemma 1, we can obtain that for any $j\in\{1,2,\cdots,p\}$,
@@
where the last equality holds because $T^{-1}\sum_{t=1}^{T}\hat{r}_{t}^{-4}=|1+H_{u}|T^{-1}\sum_{t=1}^{T}{r}_{t}^{-4}=O_{p}(\zeta_{1}^{4})$ and $\|\hat{{\bf D}}^{-1/2}\hat{\boldsymbol{\theta}}\|^{2}=O_{p}(\zeta_{1}^{-2}T^{-1})$. Then, let’s focus on the third term. We have
@@
$\leq$ $\left|1+H_{u}\right|\left\|\zeta_{1}^{-1}\frac{1}{T}\sum_{t=1}^{T}\hat{r}^{-2}_{t}\|\hat{{\bf D}}^{-1/2}\boldsymbol{\varepsilon}^{\top}{\bf F}({\bf F}^{\top}{\bf F})^{-1}{\boldsymbol{f}}_{t}\|^{2}{{\bf U}}_{t}\right\|_{\infty}$
$\leq$ $\sqrt{\left(T^{-1}\sum_{t=1}^{T}\hat{r}_{t}^{-4}\|\hat{{\bf D}}^{-1/2}\boldsymbol{\varepsilon}^{\top}{\bf F}({\bf F}^{\top}{\bf F})^{-1}{\boldsymbol{f}}_{t}\|^{4}\right)\|T^{-1}\sum_{t=1}^{T}\zeta_{1}^{-2}({{\bf U}}_{t})^{2}\|_{\infty}}\left|1+H_{u}\right|$
$=$ $O_{p}(\zeta_{1}^{2})\max_{t}\|\hat{{\bf D}}^{-1/2}\boldsymbol{\varepsilon}^{\top}{\bf F}({\bf F}^{\top}{\bf F})^{-1}{\boldsymbol{f}}_{t}\|^{2}\left|1+H_{u}\right|.$
Next, we calculate $\max_{t}\|\hat{{\bf D}}^{-1/2}\boldsymbol{\varepsilon}^{\top}{\bf F}({\bf F}^{\top}{\bf F})^{-1}{\boldsymbol{f}}_{t}\|^{2}$. Let $({{\bf D}}^{-1/2}\boldsymbol{\varepsilon}^{\top})_{i}=(\varepsilon_{i1}/d_{i},\dots,\varepsilon_{iT}/d_{i})$ is the $i$-th row vector of ${{\bf D}}^{-1/2}\boldsymbol{\varepsilon}^{\top}$. Let $\boldsymbol{\Gamma}_{i}$ is the $i$-th row vector of $\mathbf{\Gamma}$. Then, due to $\max_{1\leq i\leq N}\max_{1\leq s\leq T}\|W_{i,s}\|_{\psi_{\alpha}}\leq c_{0}$ for some constant $c_{0}>0$, we have $\|\boldsymbol{\Gamma}_{i}^{\top}\boldsymbol{W}_{s}\|_{\psi_{\alpha}}\lesssim c_{0}$ by Lemma B.4 in Koike (2021). Then, according to the sub-exponential properties, we can obtain that $E(|\boldsymbol{\Gamma}_{i}^{\top}\boldsymbol{W}_{s}|^{k})\leq c_{0}^{k}(k)^{k/\alpha}\leq(c_{0}k)^{k}$ due to $1\leq\alpha\leq 2$ for any $k\leq 1$. In addition, because of Condition (C1)-(C4), we have $|V_{ts}|<K/T$ and $E(|\varepsilon_{is}V_{ts}/d_{i}|^{k})\leq(K\underline{d}^{-1}/T)^{k}E(v_{s}^{k})E(|\boldsymbol{\Gamma}_{i}^{\top}\boldsymbol{W}_{t}|^{k})\leq(K\underline{d}^{-1}/T)^{k}E(v_{s}^{k})(c_{0}k)^{k}$, where ${\bf V}_{t}=(V_{t1},\dots,V_{tT})$. By Stirling’s approximation, we can obtain that $\varepsilon_{is}V_{ts}/d_{i}$ satisfies the condition of Bernstein. Then, according to the Bernstein inequality, we have
@@
$\left\|\zeta_{1}^{-1}\frac{1}{T}\sum_{t=1}^{T}\hat{r}^{-2}_{t}\hat{\boldsymbol{\theta}}^{\top}\hat{{\bf D}}^{-1}\boldsymbol{\varepsilon}^{\top}{\bf F}({\bf F}^{\top}{\bf F})^{-1}{\boldsymbol{f}}_{t}\hat{{\bf U}}_{t}\right\|_{\infty}=O_{p}\{T^{-1}\log(NT)\}.$
@@
$\leq$ $TE(r_{t}^{-4})/(\epsilon^{4}\zeta_{1}^{4}T)\leq C\epsilon^{-4},$
where $C$ is some positive constant. Thus, $\max_{t}\hat{r}_{t}^{-2}=O_{p}(\zeta_{1}^{2}T^{1/2})$. Then, we can derive that $\max_{t}\delta_{1t}=O_{p}(T^{-1/2})$ and $\max_{t}\delta_{1t}=O_{p}(T^{-1/4})$. Hence, we have $\left\|\zeta_{1}^{-1}T^{-1}\sum_{t=1}^{T}\delta_{1t}\hat{{\bf U}}_{t}\right\|_{\infty}=O_{p}\left(T^{-1}\right)$. Next, we are concerned about the sixth term. Similar to proof of Lemma 8 in Liu et al. (2024), we have
@@
$\leqslant$ $\zeta_{1}^{-1}\|\frac{1}{T}\sum_{t=1}^{T}\hat{r}_{t}^{-1}\hat{{\bf U}}_{t}\hat{{\bf U}}_{t}^{\top}\|_{1}\left\|\hat{{\bf D}}^{-1/2}\hat{\boldsymbol{\theta}}\right\|_{\infty}$
$\lesssim$ $N^{-1}\|\boldsymbol{{\bf R}}\|_{1}\left\|\hat{{\bf D}}^{-1/2}\hat{\boldsymbol{\theta}}\right\|_{\infty}$
$+$ $O_{p}\left(T^{-1/2}+N^{-1/6}+N^{-\delta/2}+\ T^{-1/2}(\log N)^{1/2}\left(N^{-3/2}+N^{-\delta/2}\right)\right)\left\|\hat{{\bf D}}^{-1/2}\hat{\boldsymbol{\theta}}\right\|_{\infty}$
$\quad O_{p}\left(T^{-1/2}+N^{-1/6}+N^{-\delta/2}+T^{-1/2}(\log N)^{1/2}N^{-3/2}\right)\left\|\hat{{\bf D}}^{-1/2}\hat{\boldsymbol{\theta}}\right\|_{\infty}.$
@@
$\leq$ $\sqrt{\frac{1}{T}\sum_{t=1}^{T}\hat{r}_{t}^{-2}(\hat{{\bf U}}_{t}^{\top}\hat{{\bf D}}^{-1/2}\boldsymbol{\varepsilon}^{\top}\boldsymbol{V}_{t})^{2}}\times\sqrt{\|\frac{1}{T}\sum_{t=1}^{T}\zeta_{1}^{-2}({\bf U}_{t})^{2}\|_{\infty}}$
$=$ $\sqrt{\max_{t}(\hat{{\bf U}}_{t}^{\top}\hat{{\bf D}}^{-1/2}\boldsymbol{\varepsilon}^{\top}\boldsymbol{V}_{t})^{2}}\sqrt{\frac{1}{T}\sum_{t=1}^{T}\hat{r}_{t}^{-2}}$
$\leq$ $\sqrt{\max_{t}\|\hat{{\bf D}}^{-1/2}\boldsymbol{\varepsilon}^{\top}\boldsymbol{V}_{t}\|^{2}}O_{p}(\zeta_{1})$
$=$ $O_{p}\{T^{-1/2}\log^{1/2}(NT)\}.$
@@
$\zeta_{1}^{-1}\left\|\frac{1}{T}\sum_{t=1}^{T}(1+\delta_{1t}+\delta_{2t})\hat{r}_{t}^{-1}\hat{{\bf D}}^{-1/2}\boldsymbol{\varepsilon}^{\top}\boldsymbol{V}_{t}\right\|_{\infty}$
@@
$=$ $=O_{p}\{T^{-1/2}\log^{1/2}(NT)\}.$
@@
$=$ $O_{p}\Big{\{}T^{-1}T^{-1/2}\log^{1/2}(NT)+N^{-(1/6\wedge\delta/2)}T^{-1/2}\log^{1/2}(NT)$
$+T^{-1}(\log N)^{1/2}N^{-3/2}T^{-1/2}\log^{1/2}(NT)\Big{\}}.$
@@
$T^{1/2}\hat{\mathbf{D}}^{-1/2}\hat{\boldsymbol{\theta}}=T^{-1/2}\zeta_{1}^{-1}\sum_{t=1}^{T}{\bf U}_{t}+C_{T}$ (7)

### v2

@@ A.1 Proof of Theorem 1
$-\hat{r}_{t}^{-2}\hat{\boldsymbol{\theta}}^{\top}\hat{{\bf D}}^{-1}\boldsymbol{\varepsilon}^{\top}{\bf F}({\bf F}^{\top}{\bf F})^{-1}{\boldsymbol{f}}_{t}+\delta_{1t})\hat{{\bf U}}_{t}+\frac{1}{T}\sum_{t=1}^{T}(\hat{r}_{t}^{-1}\hat{{\bf U}}_{t}\hat{{\bf U}}_{t}^{\top}\hat{{\bf D}}^{-1/2}\hat{\boldsymbol{\theta}}+\hat{r}_{t}^{-1}\hat{{\bf U}}_{t}\hat{{\bf U}}_{t}^{\top}\hat{{\bf D}}^{-1/2}\boldsymbol{\varepsilon}^{\top}{\boldsymbol{V}}_{t})$
@@
$+\hat{r}^{-1}_{t}\hat{{\bf U}}^{\top}_{t}\hat{{\bf D}}^{1/2}\boldsymbol{\varepsilon}^{\top}{\bf F}({\bf F}^{\top}{\bf F})^{-1}{\boldsymbol{f}}_{t}-\hat{r}^{-2}_{t}\hat{\boldsymbol{\theta}}^{\top}\hat{{\bf D}}^{-1}\boldsymbol{\varepsilon}^{\top}{\bf F}({\bf F}^{\top}{\bf F})^{-1}{\boldsymbol{f}}_{t}$
@@
Firstly, we focus on the first term on the left side of (2). Due to Lemma 1 and Equation (7.6) in Liu et al. (2024), we can obtain that for any $j\in\{1,2,\cdots,p\}$,
@@
where the last equality holds because $T^{-1}\sum_{t=1}^{T}\hat{r}_{t}^{-4}=\{1+o_{p}(1)\}T^{-1}\sum_{t=1}^{T}{r}_{t}^{-4}=O_{p}(\zeta_{1}^{4})$ and $\|\hat{{\bf D}}^{-1/2}\hat{\boldsymbol{\theta}}\|^{2}=O_{p}(\zeta_{1}^{-2}T^{-1})$. Then, let’s focus on the third term. We have
@@
$\leq$ $\{1+o_{p}(1)\}\left\|\zeta_{1}^{-1}\frac{1}{T}\sum_{t=1}^{T}\hat{r}^{-2}_{t}\|\hat{{\bf D}}^{-1/2}\boldsymbol{\varepsilon}^{\top}{\bf F}({\bf F}^{\top}{\bf F})^{-1}{\boldsymbol{f}}_{t}\|^{2}{{\bf U}}_{t}\right\|_{\infty}$
$\leq$ $\sqrt{\left(T^{-1}\sum_{t=1}^{T}\hat{r}_{t}^{-4}\|\hat{{\bf D}}^{-1/2}\boldsymbol{\varepsilon}^{\top}{\bf F}({\bf F}^{\top}{\bf F})^{-1}{\boldsymbol{f}}_{t}\|^{4}\right)\|T^{-1}\sum_{t=1}^{T}\zeta_{1}^{-2}({{\bf U}}_{t})^{2}\|_{\infty}}\{1+o_{p}(1)\}$
$=$ $O_{p}(\zeta_{1}^{2})\max_{t}\|\hat{{\bf D}}^{-1/2}\boldsymbol{\varepsilon}^{\top}{\bf F}({\bf F}^{\top}{\bf F})^{-1}{\boldsymbol{f}}_{t}\|^{2}\{1+o_{p}(1)\}.$
Next, we calculate $\max_{t}\|\hat{{\bf D}}^{-1/2}\boldsymbol{\varepsilon}^{\top}{\bf F}({\bf F}^{\top}{\bf F})^{-1}{\boldsymbol{f}}_{t}\|^{2}$. Let $({{\bf D}}^{-1/2}\boldsymbol{\varepsilon}^{\top})_{i}=(\varepsilon_{i1}/d_{i},\dots,\varepsilon_{iT}/d_{i})$ is the $i$-th row vector of ${{\bf D}}^{-1/2}\boldsymbol{\varepsilon}^{\top}$. Let $\boldsymbol{\Gamma}_{i}$ is the $i$-th row vector of $\mathbf{\Gamma}$. Then, due to $\max_{1\leq i\leq N}\max_{1\leq s\leq T}\|W_{i,s}\|_{\psi_{\alpha}}\leq c_{0}$ for some constant $c_{0}>0$, we have $\|\boldsymbol{\Gamma}_{i}^{\top}\boldsymbol{W}_{s}\|_{\psi_{\alpha}}\lesssim c_{0}$ by Lemma B.4 in Koike (2021). Then, according to the sub-exponential properties, we can obtain that $E(|\boldsymbol{\Gamma}_{i}^{\top}\boldsymbol{W}_{s}|^{k})\leq c_{0}^{k}(k)^{k/\alpha}\leq(c_{0}k)^{k}$ due to $1\leq\alpha\leq 2$ for any $k\leq 1$. In addition, because of Conditions (C1)-(C4), we have $|V_{ts}|<K/T$ and $E(|\varepsilon_{is}V_{ts}/d_{i}|^{k})\leq(K\underline{d}^{-1}/T)^{k}E(v_{s}^{k})E(|\boldsymbol{\Gamma}_{i}^{\top}\boldsymbol{W}_{t}|^{k})\leq(K\underline{d}^{-1}/T)^{k}E(v_{s}^{k})(c_{0}k)^{k}$, where ${\bf V}_{t}=(V_{t1},\dots,V_{tT})$. By Stirling’s approximation, we can obtain that $\varepsilon_{is}V_{ts}/d_{i}$ satisfies the condition of Bernstein. Then, according to the Bernstein inequality, we have
@@
$\left\|\zeta_{1}^{-1}\frac{1}{T}\sum_{t=1}^{T}\hat{r}^{-1}_{t}\hat{\boldsymbol{\theta}}^{\top}\hat{{\bf D}}^{-1}\boldsymbol{\varepsilon}^{\top}{\bf F}({\bf F}^{\top}{\bf F})^{-1}{\boldsymbol{f}}_{t}\hat{{\bf U}}_{t}\right\|_{\infty}=O_{p}\{T^{-1}\log(NT)\}.$
@@
$\leq$ $TE(r_{t}^{-4})/(\epsilon^{4}\zeta_{1}^{4}T)\leq C\epsilon^{-4},$ (4)
where $C$ is some positive constant. Thus, $\max_{t}\hat{r}_{t}^{-2}=O_{p}(\zeta_{1}^{2}T^{1/2})$. Then, we can derive that $\max_{t}\delta_{1t}=O_{p}(T^{-1/2})$ and $\max_{t}\delta_{2t}=O_{p}(T^{-1/4})$. Hence, we have $\left\|\zeta_{1}^{-1}T^{-1}\sum_{t=1}^{T}\delta_{1t}\hat{{\bf U}}_{t}\right\|_{\infty}=O_{p}\left(T^{-1}\right)$. Next, we are concerned about the sixth term. Similar to proof of Lemma 8 in Liu et al. (2024), we have
@@
$=$ $\left\|\zeta_{1}^{-1}\frac{1}{T}\sum_{t=1}^{T}{r}_{t}^{-1}{{\bf U}}_{t}{{\bf U}}_{t}^{\top}\hat{{\bf D}}^{-1/2}\hat{\boldsymbol{\theta}}\right\|_{\infty}\{1+o(1)\}$
$\leqslant$ $\zeta_{1}^{-1}\|\frac{1}{T}\sum_{t=1}^{T}{r}_{t}^{-1}{{\bf U}}_{t}t{{\bf U}}_{t}^{\top}\|_{1}\left\|\hat{{\bf D}}^{-1/2}\hat{\boldsymbol{\theta}}\right\|_{\infty}\{1+o(1)\}$
$\lesssim$ $N^{-1}\|\boldsymbol{{\bf R}}\|_{1}\left\|\hat{{\bf D}}^{-1/2}\hat{\boldsymbol{\theta}}\right\|_{\infty}\{1+o(1)\}$
$+$ $O_{p}\left(T^{-1/2}+N^{-1/6}+N^{-\delta/2}+\ T^{-1/2}(\log N)^{1/2}\left(N^{-3/2}+N^{-\delta/2}\right)\right)\left\|\hat{{\bf D}}^{-1/2}\hat{\boldsymbol{\theta}}\right\|_{\infty}\{1+o(1)\}$
$+$ $O_{p}\left(T^{-1/2}+N^{-1/6}+N^{-\delta/2}+T^{-1/2}(\log N)^{1/2}N^{-3/2}\right)\left\|\hat{{\bf D}}^{-1/2}\hat{\boldsymbol{\theta}}\right\|_{\infty}\{1+o(1)\}.$
@@
$=$ $\max_{1\leq i\leq N}\Big{|}\zeta_{1}^{-1}\frac{1}{T}\sum_{t=1}^{T}r_{t}^{-1}\sum_{j=1}^{N}\sum_{s=1}^{T}r_{s}U_{s,j}U_{t,j}U_{t,i}V_{ts}\Big{|}\{1+o(1)\}$
$\leq$ $\max_{1\leq i\leq N}\sqrt{\zeta_{1}^{-2}\frac{1}{T}\sum_{t=1}^{T}U_{t,i}^{2}}\sqrt{\frac{1}{T}\sum_{t=1}^{T}r_{t}^{-2}\Big{(}\sum_{j=1}^{N}\sum_{s=1}^{T}r_{s}U_{s,j}U_{t,j}V_{ts}\Big{)}^{2}}\{1+o(1)\}$
$=$ $O_{p}(1)\sqrt{\frac{1}{T}\sum_{t=1}^{T}r_{t}^{-2}\Big{(}\sum_{j=1}^{N}\sum_{s=1}^{T}r_{s}U_{s,j}U_{t,j}V_{ts}\Big{)}^{2}}.$
Because
$E\frac{1}{T}\sum_{t=1}^{T}r_{t}^{-2}\Big{(}\sum_{j=1}^{N}\sum_{s=1}^{T}r_{s}U_{s,j}U_{t,j}V_{ts}\Big{)}^{2}$
$=$ $E\frac{1}{T}\sum_{t=1}^{T}r_{t}^{-2}\sum_{j_{1}=1}^{N}\sum_{s_{1}=1}^{T}\sum_{j_{2}=1}^{N}\sum_{s_{2}=1}^{T}r_{s_{1}}U_{s_{1},j_{1}}U_{t,j_{1}}V_{ts_{1}}r_{s_{2}}U_{s_{2},j_{2}}U_{t,j_{2}}V_{ts_{2}}$
$=$ $E\frac{1}{T}\sum_{t=1}^{T}r_{t}^{-2}\sum_{j_{1}=1}^{N}\sum_{j_{2}=1}^{N}r_{t}U_{t,j_{1}}U_{t,j_{1}}V_{tt}r_{t}U_{t,j_{2}}U_{t,j_{2}}V_{tt}$
$+$ $E\frac{1}{T}\sum_{t=1}^{T}r_{t}^{-2}\sum_{j_{1}=1}^{N}\sum_{s=1,s\neq t}^{T}\sum_{j_{2}=1}^{N}r_{s}U_{s,j_{1}}U_{t,j_{1}}V_{ts}r_{s}U_{s,j_{2}}U_{t,j_{2}}V_{ts}$
$=$ $O(T^{-2}+N^{-\delta}T^{-1}),$
where the last inequality holds because $\max_{j=1,\cdots,N}\sum_{\ell=1}^{N}\left|\sigma_{j\ell}\right|\leqslant a_{0}(N)=N^{1-\delta}$, $E(U_{t,i}U_{t,j})=N^{-1}\sigma_{i,j}+O(N^{-1-\delta/2})$ by Lemma A.4(iii) in Cheng et al. (2023). So, we can conclude that $\|\zeta_{1}^{-1}\frac{1}{T}\sum_{t=1}^{T}\hat{r}_{t}^{-1}\hat{{\bf U}}_{t}\hat{{\bf U}}_{t}^{\top}\hat{{\bf D}}^{-1/2}\boldsymbol{\varepsilon}^{\top}\boldsymbol{V}_{t}\|_{\infty}=O_{p}(\max\{N^{-\delta/2}T^{-1/2},T^{-1}\})$.
@@
$\zeta_{1}^{-1}\left\|\frac{1}{T}\sum_{t=1}^{T}\hat{r}_{t}^{-1}\hat{{\bf D}}^{-1/2}\boldsymbol{\varepsilon}^{\top}\boldsymbol{V}_{t}\right\|_{\infty}$
@@
$=$ $O_{p}\{T^{-1/2}\log^{1/2}(NT)\},$
and
$\zeta_{1}^{-1}\left\|\frac{1}{T}\sum_{t=1}^{T}(\delta_{1t}+\delta_{2t})\hat{r}_{t}^{-1}\hat{{\bf D}}^{-1/2}\boldsymbol{\varepsilon}^{\top}\boldsymbol{V}_{t}\right\|_{\infty}$
$=$ $\zeta_{1}^{-1}\left\|\frac{1}{T}\sum_{t=1}^{T}{r}_{t}^{-1}(\delta_{1t}+\delta_{2t}){{\bf D}}^{-1/2}\boldsymbol{\varepsilon}^{\top}{\boldsymbol{V}}_{t}\right\|_{\infty}\{1+O_{p}(\sqrt{\log N/T})\}$
$=$ $\zeta_{1}^{-1}\max_{1\leq i\leq N}\left|T^{-1}\sum_{t=1}^{T}r_{t}^{-1}(\delta_{1t}+\delta_{2t})({{\bf D}}^{-1/2}\boldsymbol{\varepsilon}^{\top})_{i}^{\top}{\boldsymbol{V}}_{t}\right|\{1+O_{p}(\sqrt{\log N/T})\}$
$\leq$ $\zeta_{1}^{-1}\max_{1\leq t\leq T}|\delta_{1t}+\delta_{2t}|\max_{1\leq i\leq N}\max_{1\leq t\leq T}\left|({{\bf D}}^{-1/2}\boldsymbol{\varepsilon}^{\top})_{i}^{\top}{\boldsymbol{V}}_{t}\right|\left|T^{-1}\sum_{t=1}^{T}r_{t}^{-1}\right|\{1+O_{p}(\sqrt{\log N/T})\}$
$=$ $O_{p}(T^{-1/4})\max_{1\leq i\leq N}\max_{1\leq t\leq T}\left|({{\bf D}}^{-1/2}\boldsymbol{\varepsilon}^{\top})_{i}^{\top}{\boldsymbol{V}}_{t}\right|$
$=$ $O_{p}\{T^{-3/4}\log^{1/2}(NT)\}.$
@@
$=$ $O_{p}\Big{\{}T^{-1}\log^{1/2}(NT)+N^{-(1/6\wedge\delta/2)}T^{-1/2}\log^{1/2}(NT)$
$+T^{-1}(\log N)^{1/2}N^{-3/2}\log^{1/2}(NT)\Big{\}}.$
Note that
$\frac{1}{T}\sum_{t=1}^{T}{\bf U}_{t}-\frac{1}{T}r_{t}^{-1}{\bf D}^{-1/2}\boldsymbol{\varepsilon}^{\top}{\bf F}({\bf F}^{\top}{\bf F})^{-1}{\boldsymbol{f}}_{t}$
$=$ $\frac{1}{T}\sum_{t=1}^{T}(1-\sum_{s=1}^{T}r_{s}^{-1}r_{t}V_{st}){\bf U}_{t}.$
@@
$T^{1/2}\hat{\mathbf{D}}^{-1/2}\hat{\boldsymbol{\theta}}=T^{-1/2}\zeta_{1}^{-1}\sum_{t=1}^{T}(1-\sum_{s=1}^{T}r_{s}^{-1}r_{t}V_{st}){\bf U}_{t}+C_{T},$ (5)